# Colorful Pinball: Density-Weighted Quantile Regression for Conditional Guarantee of Conformal Prediction

### v2

@@ 6 Experiments
In summary, we consider eight real-world regression benchmarks spanning engineering systems, energy, transportation, and economics. Most of them are from the classic UCI repository UCIRepository. Among them, three datasets (Bike, Diamond, and Superconductivity) have one-dimensional responses, while the remaining five involve multi-dimensional responses. The covariate dimension ranges from a moderate scale (tens of features), and the sample size varies between 10k and 50k. Each dataset is randomly split into training, calibration, and test sets with an approximate ratio of 6:2:2.
@@
We consider standard baselines including Split Conformal Prediction (Split; papadopoulos2002inductive), Partition Learning Conformal Prediction (PLCP; kiyani2024conformal), Gaussian Scoring braun2025multivariate, CQR romano2019conformalized, RCP plassier2025rectifying. In addition, to separately validate the value of our approximation described in Proposition 3.4, we select a parametric likelihood on nonconformity score, the Asymmetric Laplace Distribution (ALD), of which the maximum likelihood estimation (MLE) is equivalent to minimizing scaled pinball loss and a regularization term on scales $\sigma(x)$. We construct two baselines based on CQR and RCP through substituting the quantile regression with MLE of ALD, which does not involve the rectification procedures. We provide the details of baselines in Appendix A.2.
Except for Gaussian Scoring, we adopt $l_{\infty}$ norm of residuals as nonconformity scores diquigiovanni2024importance, which reduces to the standard absolute residual score for the case of one-dimensional labels.
@@
To approximately evaluate the conditional coverage in real-world datasets, where we have no access to the data generation mechanism, we adopt clustering (K-means; braun2025conditionalcoveragediagnosticsconformal) for MSCE and random projection for Worst-Slice Coverage (WSC; cauchois2021knowing, romano2020classification). In general, given marginal validity, both lower MSCE and higher WSC imply better concentration of conditional coverage probabilities, and WSC is additionally a robustness measure that characterizes the worst-case coverage. Throughout the experiments, we set the target coverage level $\tau=90\%$. We also report the mean (per dimension) of the log volume of the predictive set to characterize the efficiency. All methods are run with 20 Monte Carlo repetitions to evaluate the metrics, and we report $\operatorname{mean}\pm\operatorname{std}$. The details of metrics are provided in Appendix A.4.
Table 1: Comparison of MSCE. Mean Squared Coverage Error (MSCE, $\downarrow$) on 8 high-dimensional benchmarks. Best results are bolded.
@@
Split 0.0030 ± 0.0007 0.0127 ± 0.0029 0.0033 ± 0.0007 0.0349 ± 0.0120 0.0035 ± 0.0009 0.0073 ± 0.0008 0.0045 ± 0.0019 0.0123 ± 0.0017
PLCP 0.0018 ± 0.0007 0.0022 ± 0.0027 0.0008 ± 0.0008 0.0172 ± 0.0100 0.0028 ± 0.0014 0.0014 ± 0.0007 0.0034 ± 0.0026 0.0016 ± 0.0012
Gaussian-Scoring 0.0010 ± 0.0005 0.0006 ± 0.0004 0.0005 ± 0.0002 0.0169 ± 0.0136 0.0021 ± 0.0028 0.0031 ± 0.0013 0.0020 ± 0.0019 0.0071 ± 0.0018
CQR 0.0015 ± 0.0010 0.0026 ± 0.0010 0.0006 ± 0.0004 0.0207 ± 0.0079 0.0013 ± 0.0005 0.0011 ± 0.0008 0.0013 ± 0.0013 0.0068 ± 0.0009
CQR-ALD 0.0012 ± 0.0010 0.0013 ± 0.0005 0.0006 ± 0.0003 0.0170 ± 0.0112 0.0010 ± 0.0005 0.0009 ± 0.0006 0.0010 ± 0.0005 0.0046 ± 0.0013
RCP 0.0016 ± 0.0005 0.0033 ± 0.0008 0.0006 ± 0.0003 0.0048 ± 0.0034 0.0013 ± 0.0003 0.0017 ± 0.0005 0.0009 ± 0.0005 0.0039 ± 0.0007
RCP-ALD 0.0010 ± 0.0004 0.0019 ± 0.0008 0.0004 ± 0.0002 0.0089 ± 0.0066 0.0011 ± 0.0004 0.0011 ± 0.0006 0.0007 ± 0.0003 0.0031 ± 0.0006
CPCP 0.0011 ± 0.0006 0.0006 ± 0.0003 0.0005 ± 0.0002 0.0020 ± 0.0011 0.0003 ± 0.0002 0.0008 ± 0.0004 0.0010 ± 0.0006 0.0015 ± 0.0005
CPCP (Clip+Mix) 0.0008 ± 0.0003 0.0004 ± 0.0002 0.0003 ± 0.0001 0.0017 ± 0.0008 0.0003 ± 0.0001 0.0007 ± 0.0004 0.0003 ± 0.0001 0.0010 ± 0.0003
Table 2: Comparison of WSC. Worst-Slice Coverage (WSC, $\uparrow$) on 8 high-dimensional benchmarks. Best results are bolded.
@@
Split 0.8083 ± 0.0265 0.6492 ± 0.0180 0.8134 ± 0.0183 0.5703 ± 0.0574 0.7470 ± 0.0097 0.7824 ± 0.0237 0.7239 ± 0.0194 0.7462 ± 0.0165
PLCP 0.8534 ± 0.0242 0.8286 ± 0.0585 0.8711 ± 0.0218 0.6821 ± 0.0925 0.7803 ± 0.0441 0.8661 ± 0.0218 0.7936 ± 0.0499 0.8630 ± 0.0198
Gaussian-Scoring 0.8605 ± 0.0319 0.8607 ± 0.0140 0.8727 ± 0.0180 0.6723 ± 0.0870 0.8251 ± 0.0530 0.8314 ± 0.0239 0.8287 ± 0.0356 0.8017 ± 0.0224
CQR 0.8653 ± 0.0296 0.8162 ± 0.0166 0.8752 ± 0.0170 0.6440 ± 0.0560 0.8284 ± 0.0268 0.8682 ± 0.0224 0.8313 ± 0.0204 0.8050 ± 0.0162
CQR-ALD 0.8699 ± 0.0265 0.8457 ± 0.0148 0.8750 ± 0.0177 0.6724 ± 0.0708 0.8462 ± 0.0212 0.8753 ± 0.0196 0.8378 ± 0.0250 0.8121 ± 0.0195
RCP 0.8662 ± 0.0285 0.7952 ± 0.0155 0.8780 ± 0.0198 0.7963 ± 0.0546 0.8291 ± 0.0155 0.8572 ± 0.0223 0.8272 ± 0.0168 0.8318 ± 0.0117
RCP-ALD 0.8722 ± 0.0210 0.8310 ± 0.0173 0.8860 ± 0.0125 0.7459 ± 0.0827 0.8360 ± 0.0164 0.8709 ± 0.0217 0.8570 ± 0.0142 0.8428 ± 0.0139
CPCP 0.8757 ± 0.0189 0.8685 ± 0.0173 0.8824 ± 0.0180 0.8323 ± 0.0374 0.8886 ± 0.0145 0.8618 ± 0.0205 0.8599 ± 0.0135 0.8518 ± 0.0192
CPCP (Clip+Mix) 0.8932 ± 0.0142 0.8772 ± 0.0136 0.8857 ± 0.0126 0.8346 ± 0.0345 0.8833 ± 0.0138 0.8790 ± 0.0196 0.8805 ± 0.0124 0.8692 ± 0.0143
@@ 6.1 Results analysis
We present the main part of results in Table 1 and 2, and all results with ablation studies can be found in Appendix A.5. In summary, our methodology achieves impressive improvement on conditional coverage across all benchmarks, with a medium predictive set size among all baselines (per set size, see Table 6). Specifically, the improvement on WSC is remarkable, which would be more pronounced if considering $(\tau-\mathrm{WSC})$, i.e, the worst-case miscoverage probability. In addition, we observe that the advantage of our methods becomes more prominent when standard Split CP performs badly, e.g., on Diamond and Naval, which corresponds to a more challenging setting. Moreover, we find that the parametric likelihood (ALD) indeed brings benefits in most cases, which further validates our approximation. Finally, our two additional mechanisms for stability work very well, as also further demonstrated in Table 4 and 5.
Overall, these results empirically corroborate our claim: although density-weighted optimization introduces additional estimation error, it consistently yields superior conditional coverage in the finite-sample regime, where objective misalignment dominates asymptotic considerations.
@@ A.1 Dataset description
We introduce our benchmarks in alphabetical order, following the sequence presented in our tables. First, we provide a summary of all eight datasets, after which we detail each one individually. We note that the majority of these datasets are sourced from the UCI repository UCIRepository.
@@ A.2 Baseline details
Partition Learning Conformal Prediction (PLCP; kiyani2024conformal) aims to improve upon the marginal nature of Split CP by partitioning the feature space into learned regions. It optimizes these partitions to group samples of similar difficulty and applies Split CP locally within each group. This approach is formulated as a co-training procedure:
@@
where $G$ denotes the number of distinct groups (we use $G$ to avoid conflict with the calibration sample size $m$, differing from the notation in kiyani2024conformal), and $h:\mathcal{X}\rightarrow\Delta_{G}$ is a partition function assigning samples to groups.
@@
As described in braun2025multivariate, this method assumes a multivariate Gaussian density for $Y\mid X$. On the training set, we train two MLPs to simultaneously predict the conditional mean $\hat{\mu}(X)$ and covariance matrix $\hat{\Sigma}(X)$ by minimizing the negative log-likelihood (NLL). The nonconformity score is defined as the Mahalanobis distance:
@@
which reduces to the standard normalized residual $S(X,Y)=|Y-\hat{\mu}(X)|/\hat{\sigma}(X)$ lei2018distribution when $Y$ is one-dimensional.
@@
Conformalized Quantile Regression (CQR; romano2019conformalized) constructs adaptive intervals by directly estimating conditional quantiles on the training set, rather than estimating the conditional mean $\hat{\mu}$. We train two MLPs to output the lower ($\hat{q}_{\alpha/2}$) and upper ($\hat{q}_{1-\alpha/2}$) quantiles by minimizing the pinball loss. The nonconformity score is defined as:
@@ A.4 Metric details
For MSCE, we use K-means to partition samples into $K$ groups. There is a fundamental trade-off here: increasing $K$ approximates the conditional coverage with finer granularity, but reduces the sample size within each group, thereby decreasing the precision of the empirical coverage estimate. Our guiding principle is to ensure sufficient samples (on the order of hundreds) within each group to evaluate coverage with percent-level precision. Consequently, we examine two settings: $K=10$ and $K=50$, as reported in Tables 4 and 7.
@@
For WSC, we adopt the evaluation framework of romano2020classification, cauchois2021knowing. We first define a slab in the covariate space as:

### v3

@@ 6 Experiments
We consider eight real-world regression benchmarks spanning engineering systems, energy, transportation, and economics. Among them, three datasets (Bike, Diamond, and Superconductivity) have one-dimensional responses, while the remaining five involve multi-dimensional responses. The covariate dimension ranges from a moderate scale (tens of features), and the sample size varies between 10k and 50k. Each dataset is randomly split into training, calibration, and test sets with a ratio of 6:2:2.
@@
We consider standard baselines including Split Conformal Prediction (Split; Papadopoulos et al., 2002), Partition Learning Conformal Prediction (PLCP; Kiyani et al., 2024), Gaussian Scoring Braun et al. [2025a], CQR Romano et al. [2019], RCP Plassier et al. [2025a]. In addition, to separately validate the value of our approximation described in Proposition 3.4, we select a parametric likelihood on nonconformity score, the Asymmetric Laplace Distribution (ALD), of which the maximum likelihood estimation (MLE) is equivalent to minimizing scaled pinball loss and a regularization term on scales $\sigma(x)$. We construct two baselines based on CQR and RCP through substituting the quantile regression with MLE of ALD. We provide the details of baselines in Appendix A.2.
Except for Gaussian Scoring, we adopt $l_{\infty}$ norm of residuals as nonconformity scores Diquigiovanni et al. [2024], which reduces to the standard absolute residual score for the case of one-dimensional labels.
@@
To evaluate the conditional coverage in real-world datasets, where we have no access to the data generation mechanism, we adopt MSCE (approximated by K-means), Worst-Slice Coverage (WSC; Cauchois et al., 2021, Romano et al., 2020), excess risk of the target coverage ($\ell$-ERT; Braun et al., 2025b). In general, given marginal validity, better concentration of conditional coverage probabilities is indicated by lower MSCE, higher WSC, and lower $\ell$-ERT. Moreover, WSC also serves as a robustness measure, characterizing the worst-case coverage. Throughout the experiments, we set the target coverage level $\tau=90\%$. We also report the mean (per dimension) of the log volume of the predictive set to characterize the efficiency. All methods are run with 20 Monte Carlo repetitions for evaluation, and we report $\operatorname{mean}\pm\operatorname{std}$. Metric details and related results are provided in Appendices A.4 and A.5, respectively.
Table 1: Comparison of Mean Squared Coverage Error (MSCE, $\downarrow$). Best results are bolded.
@@
Split 0.0031 ± 0.0008 0.0118 ± 0.0035 0.0033 ± 0.0006 0.0351 ± 0.0064 0.0039 ± 0.0010 0.0082 ± 0.0010 0.0125 ± 0.0031 0.0123 ± 0.0019
PLCP 0.0021 ± 0.0011 0.0032 ± 0.0033 0.0008 ± 0.0009 0.0162 ± 0.0087 0.0026 ± 0.0013 0.0017 ± 0.0008 0.0027 ± 0.0038 0.0031 ± 0.0033
Gaussian-Scoring 0.0011 ± 0.0004 0.0006 ± 0.0003 0.0004 ± 0.0002 0.0129 ± 0.0095 0.0041 ± 0.0041 0.0026 ± 0.0011 0.0016 ± 0.0014 0.0057 ± 0.0026
CQR 0.0011 ± 0.0007 0.0010 ± 0.0004 0.0006 ± 0.0002 0.0120 ± 0.0064 0.0012 ± 0.0005 0.0009 ± 0.0005 0.0016 ± 0.0009 0.0061 ± 0.0007
CQR-ALD 0.0008 ± 0.0003 0.0007 ± 0.0007 0.0006 ± 0.0003 0.0055 ± 0.0059 0.0009 ± 0.0006 0.0015 ± 0.0010 0.0021 ± 0.0029 0.0045 ± 0.0009
RCP 0.0010 ± 0.0005 0.0013 ± 0.0006 0.0006 ± 0.0004 0.0029 ± 0.0020 0.0007 ± 0.0003 0.0013 ± 0.0011 0.0009 ± 0.0003 0.0030 ± 0.0005
RCP-ALD 0.0013 ± 0.0010 0.0010 ± 0.0006 0.0005 ± 0.0003 0.0080 ± 0.0102 0.0009 ± 0.0007 0.0011 ± 0.0005 0.0010 ± 0.0007 0.0025 ± 0.0008
CPCP 0.0009 ± 0.0004 0.0009 ± 0.0004 0.0004 ± 0.0002 0.0019 ± 0.0010 0.0003 ± 0.0002 0.0014 ± 0.0011 0.0009 ± 0.0006 0.0025 ± 0.0009
CPCP (Clip+Mix) 0.0008 ± 0.0002 0.0004 ± 0.0002 0.0004 ± 0.0002 0.0019 ± 0.0009 0.0003 ± 0.0001 0.0007 ± 0.0004 0.0004 ± 0.0002 0.0012 ± 0.0004
Table 2: Comparison of Worst-Slice Coverage (WSC, $\uparrow$). Best results are bolded.
@@
Split 0.8133 ± 0.0277 0.6480 ± 0.0206 0.8192 ± 0.0159 0.5428 ± 0.0446 0.7435 ± 0.0156 0.7712 ± 0.0266 0.6797 ± 0.0234 0.7623 ± 0.0195
PLCP 0.8559 ± 0.0341 0.8068 ± 0.0680 0.8690 ± 0.0229 0.6837 ± 0.0623 0.7891 ± 0.0514 0.8529 ± 0.0268 0.8123 ± 0.0645 0.8413 ± 0.0369
Gaussian-Scoring 0.8611 ± 0.0278 0.8613 ± 0.0170 0.8804 ± 0.0166 0.6928 ± 0.0911 0.7943 ± 0.0646 0.8434 ± 0.0232 0.8429 ± 0.0355 0.8259 ± 0.0347
CQR 0.8641 ± 0.0245 0.8563 ± 0.0159 0.8801 ± 0.0217 0.6997 ± 0.0837 0.8393 ± 0.0201 0.8704 ± 0.0235 0.8175 ± 0.0253 0.8149 ± 0.0188
CQR-ALD 0.8696 ± 0.0246 0.8735 ± 0.0259 0.8807 ± 0.0156 0.7836 ± 0.0745 0.8513 ± 0.0364 0.8622 ± 0.0246 0.8363 ± 0.0227 0.8222 ± 0.0193
RCP 0.8849 ± 0.0240 0.8448 ± 0.0237 0.8752 ± 0.0158 0.8002 ± 0.0457 0.8627 ± 0.0149 0.8638 ± 0.0260 0.8515 ± 0.0196 0.8516 ± 0.0198
RCP-ALD 0.8743 ± 0.0250 0.8558 ± 0.0178 0.8872 ± 0.0174 0.7596 ± 0.0803 0.8597 ± 0.0331 0.8779 ± 0.0181 0.8561 ± 0.0226 0.8582 ± 0.0173
CPCP 0.8820 ± 0.0132 0.8589 ± 0.0216 0.8844 ± 0.0125 0.8169 ± 0.0408 0.8864 ± 0.0119 0.8733 ± 0.0251 0.8548 ± 0.0172 0.8462 ± 0.0266
CPCP (Clip+Mix) 0.8882 ± 0.0250 0.8802 ± 0.0099 0.8912 ± 0.0139 0.8320 ± 0.0304 0.8912 ± 0.0126 0.8771 ± 0.0220 0.8759 ± 0.0158 0.8715 ± 0.0141
Table 3: Comparison of $\ell_{1}$ Excess Risk of the Target coverage ($\ell_{1}$-ERT, $\downarrow$). Best results are bolded.
Split 0.0602 ± 0.0071 0.1223 ± 0.0050 0.0374 ± 0.0034 0.1536 ± 0.0092 0.1044 ± 0.0046 0.0956 ± 0.0058 0.1120 ± 0.0058 0.1079 ± 0.0031
PLCP 0.0424 ± 0.0167 0.0642 ± 0.0310 0.0198 ± 0.0075 0.0789 ± 0.0279 0.0809 ± 0.0295 0.0466 ± 0.0092 0.0573 ± 0.0262 0.0656 ± 0.0182
Gaussian-Scoring 0.0337 ± 0.0086 0.0330 ± 0.0057 0.0168 ± 0.0079 0.0826 ± 0.0327 0.0733 ± 0.0349 0.0443 ± 0.0105 0.0371 ± 0.0146 0.0996 ± 0.0089
CQR 0.0307 ± 0.0136 0.0384 ± 0.0082 0.0170 ± 0.0070 0.0783 ± 0.0282 0.0506 ± 0.0112 0.0270 ± 0.0099 0.0510 ± 0.0117 0.0762 ± 0.0035
CQR-ALD 0.0274 ± 0.0114 0.0301 ± 0.0131 0.0177 ± 0.0041 0.0476 ± 0.0268 0.0404 ± 0.0204 0.0328 ± 0.0127 0.0461 ± 0.0092 0.0776 ± 0.0051
RCP 0.0208 ± 0.0076 0.0446 ± 0.0083 0.0155 ± 0.0067 0.0353 ± 0.0184 0.0371 ± 0.0099 0.0329 ± 0.0108 0.0342 ± 0.0078 0.0574 ± 0.0052
RCP-ALD 0.0274 ± 0.0142 0.0373 ± 0.0100 0.0157 ± 0.0051 0.0557 ± 0.0343 0.0427 ± 0.0185 0.0309 ± 0.0084 0.0322 ± 0.0132 0.0495 ± 0.0059
CPCP 0.0195 ± 0.0097 0.0379 ± 0.0105 0.0129 ± 0.0047 0.0271 ± 0.0148 0.0166 ± 0.0062 0.0370 ± 0.0120 0.0295 ± 0.0076 0.0632 ± 0.0090
CPCP (Clip+Mix) 0.0178 ± 0.0076 0.0219 ± 0.0048 0.0116 ± 0.0067 0.0268 ± 0.0102 0.0155 ± 0.0061 0.0267 ± 0.0065 0.0182 ± 0.0050 0.0433 ± 0.0048
@@ 6.1 Results Analysis
We present partial results in Tables 3 and 3; complete results including ablation studies are provided in Appendix A.5. In summary, our methodology achieves significant improvement on conditional coverage across all benchmarks, with a medium predictive set size among all baselines (per set size, see Table 10). Specifically, the improvement in WSC is substantial; this advantage is even more pronounced when considering $(\tau-\mathrm{WSC})$, which represents the worst-case miscoverage probability. Moreover, we observe notable improvements in $\ell$-ERT, as shown in Tables 8 and 9. We note that $\ell$-ERT is particularly effective for high-dimensional $X$.
Regarding the ablation studies, we first observe that the parametric likelihood (ALD) yields consistent benefits, further validating the approximation in Proposition 3.4. Second, our two additional stability mechanisms demonstrate high effectiveness. Finally, as shown in Appendix A.5, the influence of $\delta$ is relatively minor (comparing our default $\delta=2\%$ against $1\%$ and $5\%$). Furthermore, the limited benefit observed from pure multitasking (co-training three quantiles) further highlights the significance of our specific methodological contributions.
@@ A.1 Dataset description
We introduce our benchmarks in alphabetical order, following the sequence presented in our tables. First, we provide a summary of all eight datasets, after which we detail each one individually. We note that the majority of these datasets are sourced from the UCI repository Kelly et al. [2025].
@@ A.2 Baseline details
Partition Learning Conformal Prediction (PLCP; Kiyani et al., 2024) aims to improve upon the marginal nature of Split CP by partitioning the feature space into learned regions. It optimizes these partitions to group samples of similar difficulty and applies Split CP locally within each group. This approach is formulated as a co-training procedure:
@@
where $G$ denotes the number of distinct groups (we use $G$ to avoid conflict with the calibration sample size $m$, differing from the notation in Kiyani et al., 2024), and $h:\mathcal{X}\rightarrow\Delta_{G}$ is a partition function assigning samples to groups.
@@
As described in Braun et al. [2025a], this method assumes a multivariate Gaussian density for $Y\mid X$. On the training set, we train two MLPs to simultaneously predict the conditional mean $\hat{\mu}(X)$ and covariance matrix $\hat{\Sigma}(X)$ by minimizing the negative log-likelihood (NLL). The nonconformity score is defined as the Mahalanobis distance:
@@
which reduces to the standard normalized residual $S(X,Y)=|Y-\hat{\mu}(X)|/\hat{\sigma}(X)$ Lei et al. [2018] when $Y$ is one-dimensional.
@@
Conformalized Quantile Regression (CQR; Romano et al., 2019) constructs adaptive intervals by directly estimating conditional quantiles on the training set, rather than estimating the conditional mean $\hat{\mu}$. We train two MLPs to output the lower ($\hat{q}_{\alpha/2}$) and upper ($\hat{q}_{1-\alpha/2}$) quantiles by minimizing the pinball loss. The nonconformity score is defined as:
@@
RCP-Multihead.
We examine this baseline for examining the impact of co-training three quantiles, i.e., the multitasking step. In this baseline, we just co-train the three quantiles and use the central quantile estimator, $\hat{q}_{\tau}$ to implement the subsequent RCP procedures.
@@ A.4 Metric details
For MSCE, we use K-means to partition samples into $K$ groups Braun et al. [2025b]. There is a fundamental trade-off here: increasing $K$ approximates the conditional coverage with finer granularity, but reduces the sample size within each group, thereby decreasing the precision of the empirical coverage estimate. Our guiding principle is to ensure sufficient samples (on the order of hundreds) within each group to evaluate coverage with percent-level precision. Consequently, we examine two settings: $K=10$ and $K=30$, as reported in Tables 5 and 6. We note that this approach depends on Euclidean distance, which becomes increasingly unreliable in high dimensions, rendering it ineffective for high-dimensional $X$.
@@
Romano et al. [2020], Cauchois et al. [2021] propose the worst-slab coverage (WSC). We first define a slab in the covariate space as:
@@
$\ell$-ERT
Braun et al. [2025b] introduce the Excess Risk of the Target coverage (ERT), which casts coverage evaluation as a binary classification problem. This approach relies on the insight that under perfect conditional coverage, the coverage indicator $Z=\mathbb{I}\{Y\in\mathcal{C}_{\alpha}(X)\}$ behaves as a Bernoulli variable with constant probability $1-\alpha$, independent of $X$. Consequently, any classifier $h(X)$ that achieves a lower risk than the constant predictor $1-\alpha$ exposes a violation of conditional validity. The metric is defined as the risk reduction:
-$\ell\text{-ERT}=\mathcal{R}_{\ell}(1-\alpha)-\mathcal{R}_{\ell}(h),$ (38)
where $\mathcal{R}_{\ell}$ is the risk under a proper loss function $\ell$. Depending on the choice of $\ell$, this metric estimates different distances; specifically, we consider $\ell_{1}$-ERT and $\ell_{2}$-ERT, which provide lower-bound estimates for $\mathbb{E}[|\pi(X)-\tau|]$ and $\mathbb{E}[(\pi(X)-\tau)^{2}]$ (i.e., MSCE), respectively. In our experiments, we employ Logistic Regression as the classifier $h$ to estimate these quantities.